# Diffusion Denoising as a Certified Defense against Clean-label Poisoning

### v1

@@ 3.3 Prediction and Certification
Now we present algorithms for running a prediction with the classifiers trained on $D_{ptr}$ and certifying the robustness of the model’s prediction on a test-time sample $x$. Our work extends the algorithms presented by Cohen et al. (2019) to clean-label poisoning settings, as shown in Algorithm 2.
Prediction. To compute the prediction for a test-time sample $x$, we make an adaptation of the standard randomized smoothing. We train $n$ classifiers on $n$ different noised-then-denoised training datasets, and return the predictions of these classifiers on the target $x$, with the algorithm DenoiseTrainClassify($\cdot$). We run this algorithm for a sufficient number of times (e.g., over 1000 times). The prediction output is then the majority-voted label from these classifiers.
Certification process exactly follows the standard randomized smoothing. We first count the number of occurrences of the most likely label $\hat{c_{A}}$ compared to any other runner-up label, and from this can derive a radius (for the training set perturbations) on which $x$ is guaranteed to be robust.
Figure 1: Certified radius and accuracy attained by denoising the CIFAR10 training data with different $\sigma$ values.
Results.
We train 10k CIFAR10 models using an efficient training pipeline that trains a single model in $15$ seconds on an A100 GPU (for a total of 1.7 GPU days of compute). These models have a clean accuracy of between $92\%$ and $58\%$ depending on the level of noise $\sigma$ introduced, and for the first time can certify non-trivial robustness to clean-label poisoning attacks. A full curve comparing certified accuracy for all given perturbation radii is shown in Figure 1.
Our defense against $\ell_{2}$ attacks at $\sigma$ (%) Our defense against $\ell_{\infty}$ attacks at $\sigma$ (%) Poisoning attacks Knowledge ${}^{\dagger}$0.0 0.1 0.25 0.5 1.0 ${}^{\dagger}$0.0 0.1 0.25 0.5 1.0 Poison Frog! WB ${}^{(93.6)}$99.0 ${}^{(93.3)}$0.0 ${}^{(91.8)}$1.0 ${}^{(84.8)}$0.0 ${}^{(79.9)}$1.0 ${}^{(93.6)}$68.8 ${}^{(93.3)}$0.0 ${}^{(92.7)}$0.0 ${}^{(90.8)}$0.0 ${}^{(87.4)}$0.0 Convex Polytope ${}^{(93.7)}$16.2 ${}^{(93.2)}$0.0 ${}^{(91.7)}$0.0 ${}^{(86.6)}$0.0 ${}^{(77.0)}$0.0 ${}^{(93.7)}$12.2 ${}^{(93.3)}$0.0 ${}^{(92.7)}$0.0 ${}^{(90.8)}$1.0 ${}^{(87.5)}$0.0 Bullseye Polytope ${}^{(93.5)}$100 ${}^{(93.3)}$4.0 ${}^{(92.6)}$0.0 ${}^{(87.5)}$0.0 ${}^{(79.2)}$1.0 ${}^{(93.5)}$100 ${}^{(93.3)}$0.0 ${}^{(92.7)}$0.0 ${}^{(90.8)}$1.0 ${}^{(87.5)}$0.0 Label-consistent Backdoor - ${}^{(93.2)}$1.0 ${}^{(93.3)}$0.0 ${}^{(92.6)}$0.0 ${}^{(90.8)}$1.0 ${}^{(87.5)}$0.0 Hidden Trigger Backdoor - ${}^{(93.4)}$7.0 ${}^{(93.3)}$0.0 ${}^{(92.6)}$0.0 ${}^{(90.8)}$0.0 ${}^{(87.5)}$0.0 Poison Frog! BB ${}^{(91.6)}$10.0 ${}^{(91.2)}$0.0 ${}^{(89.6)}$0.5 ${}^{(82.9)}$0.0 ${}^{(77.8)}$2.0 ${}^{(91.7)}$2.5 ${}^{(91.3)}$0.0 ${}^{(90.3)}$0.0 ${}^{(88.8)}$0.5 ${}^{(86.2)}$1.0 Convex Polytope ${}^{(91.7)}$3.0 ${}^{(91.0)}$0.0 ${}^{(89.5)}$0.0 ${}^{(84.6)}$0.5 ${}^{(73.6)}$1.0 ${}^{(91.8)}$2.5 ${}^{(91.3)}$0.0 ${}^{(90.3)}$0.0 ${}^{(88.8)}$0.5 ${}^{(86.2)}$1.0 Bullseye Polytope ${}^{(91.6)}$9.0 ${}^{(91.3)}$0.0 ${}^{(90.3)}$0.0 ${}^{(85.5)}$0.0 ${}^{(76.3)}$1.0 ${}^{(91.6)}$8.0 ${}^{(91.3)}$0.0 ${}^{(90.3)}$0.0 ${}^{(88.8)}$0.5 ${}^{(86.2)}$0.5 Label-consistent Backdoor - ${}^{(91.5)}$1.0 ${}^{(91.3)}$0.0 ${}^{(90.3)}$0.0 ${}^{(88.8)}$0.0 ${}^{(86.2)}$1.5 Hidden Trigger Backdoor - ${}^{(91.6)}$4.0 ${}^{(91.2)}$1.0 ${}^{(90.3)}$1.0 ${}^{(89.3)}$1.5 ${}^{(86.3)}$1.5
Table 1: Defense effectiveness in transfer-learning scenarios (CIFAR10). We measure the clean accuracy and attack success of the models trained on the denoised training set. Each cell shows the accuracy in the parentheses and the attack success outside. Note that ${}^{\dagger}$ indicates the runs with $\sigma$=0.0, the same as our baseline that trains models without any denoising.
@@ 4 Empirical Evaluation
We empirically evaluate the effectiveness of our certified defense against clean-label poisoning. We adapt the poisoning benchmark developed by Schwarzschild et al. (2021). This benchmark runs clean-label poisoning attacks with the same attack configurations, allowing us to compare our defense’s effectiveness across many existing attacks.
Poisoning attacks. We evaluate our defense against seven clean-label poisoning attacks: four targeted poisoning (Poison Frogs (Shafahi et al., 2018), Convex Polytope (Zhu et al., 2019), Bullseye Polytope (Aghakhani et al., 2021), Witches’ Brew (Geiping et al., 2021b)) and three backdoor attacks (Label-consistent Backdoor (Turner et al., 2019), Hidden-Trigger Backdoor (Saha et al., 2020), and Sleeper Agent (Souri et al., 2022)). The first three attacks operate in the transfer-learning scenarios (i.e., they assume that the victim fine-tunes a model from some given initialization), whereas the rest assumes the victim trains a model from-scratch. While our defense offers provable guarantees only to targeted poisoning attacks, we test if the defense can mitigate clean-label backdooring like the prior work did.
@@
Attack scenarios. We evaluate against both the transfer-learning and training from-scratch attacks. In transfer-learning, we fine-tune a pre-trained model on the tampered training set while we train a randomly initialized model in the training from-scratch scenario. We also evaluate against the attackers with two types of knowledge: white-box and black-box. The white-box attackers know the victim model’s architecture (and its parameters), while the black-box attackers do not. More details are in Appendix A.
Our defense against $\ell_{2}$ attacks at $\sigma$ (%) Our defense against $\ell_{\infty}$ attacks at $\sigma$ (%) Poisoning attacks Knowledge ${}^{\dagger}$0.0 0.1 0.25 0.5 1.0 ${}^{\dagger}$0.0 0.1 0.25 0.5 1.0 Witches’ Brew WB ${}^{(92.2)}$71.0 ${}^{(86.4)}$54.0 ${}^{(72.3)}$10.0 ${}^{(46.7)}$11.0 ${}^{(42.5)}$10.0 ${}^{(92.3)}$65.0 ${}^{(86.5)}$9.0 ${}^{(71.9)}$3.0 ${}^{(46.0)}$9.0 ${}^{(41.3)}$7.0 Sleeper Agent Backdoor ${}^{(92.4)}$40.5 ${}^{(84.4)}$66.5 ${}^{(71.8)}$19.5 ${}^{(46.8)}$13.0 ${}^{(39.9)}$10.5 ${}^{(92.4)}$35.0 ${}^{(86.1)}$17.0 ${}^{(73.0)}$8.0 ${}^{(47.0)}$9.5 ${}^{(39.7)}$10.0 Witches’ Brew BB ${}^{(90.1)}$45.5 ${}^{(85.9)}$28.0 ${}^{(75.5)}$4.0 ${}^{(58.8)}$7.0 ${}^{(49.0)}$10.0 ${}^{(90.0)}$33.5 ${}^{(85.8)}$3.5 ${}^{(75.5)}$2.5 ${}^{(58.8)}$6.0 ${}^{(48.7)}$6.5 Sleeper Agent Backdoor ${}^{(90.0)}$39.5 ${}^{(85.0)}$44.5 ${}^{(75.1)}$14.5 ${}^{(58.6)}$9.5 ${}^{(49.0)}$8.0 ${}^{(90.0)}$18.5 ${}^{(85.6)}$11.5 ${}^{(75.5)}$7.0 ${}^{(58.8)}$8.0 ${}^{(48.4)}$8.0
Table 2: Defense effectiveness in training from-scratch scenarios (CIFAR10). We measure the accuracy and attack success of the models trained on the denoised training set. Each cell shows the accuracy in the parentheses and the attack success outside. Note that ${}^{\dagger}$ indicates the runs with $\sigma$=0.0, the same as our baseline that trains models without any denoising. We use an ensemble of four models, and WB and BB stand for the white-box and the black-box attacks, respectively.
@@ 4.1 Effectiveness of Our Denoising Defense
We first show to what extent our defense makes the seven poisoning attacks ineffective. Our defense is attack-agnostic. A defender does not know whether the training data contains poisoning samples or not; instead, the defender always trains a model on the denoised training set. We measure the trained model’s accuracy and attack success and compare them with the baseline, where a model is trained directly without denoising. We evaluate with different defense strengths: $\sigma\in\{0.1,0.25,0.5,1.0\}$. The attacks are with the perturbation bound of 16-pixels in $\ell_{\infty}$-norm and that of 3.5–6.5 in $\ell_{2}$-norm. We extended many poisoning attacks for supporting $\ell_{2}$ bounds. Both $\ell_{\infty}$ and $\ell_{2}$ bounds we use are comparable to each other. We show the conversion between the two bounds in Appendix B. Due to the page limit, we show the results from Tiny ImageNet and the results from attacking with different $\ell_{p}$-norms in Appendix C.
@@
Transfer-learning scenarios. Table 1 summarizes our results. Our defense is also effective in mitigating the attacks in transfer-learning. Each cell contains the accuracy and the attack success rate of a model trained on the denoised training set. All the numbers are the averaged values over 100 attacks. Against the white-box attacks, we reduce the success rate of the targeted poisoning attacks to 0–4% at $\sigma$ of 0.1. In the black-box attacks, the targeted poisoning is less effective (with 3–10% success), but still, the defense reduces the success to 0–1%. We observe that the two clean-label backdoor attacks show 1–7% success regardless of the attacker’s knowledge. We thus exclude these attacks from the subsequent evaluation. Using $\sigma$ greater than 0.5 increases the attack success by 0.5–2%. But, this may not because of the successful poisoning attacks but due to the significant decrease in a model’s accuracy (3-9%). Increasing $\sigma$ may defeat a stronger attack but significantly decrease a model’s accuracy. This result is interesting, as in our theoretical analysis shown in Sec 3, our defense needs $\sigma$ of 0.25, while 0.1 is sufficient to defeat both the attacks.
@@
Training from-scratch scenarios. Table 2 shows our results with the same format as in Table 1. We demonstrate that our defense mitigates both attacks at $\sigma=$ 0.25, aligned to what we’ve seen in our theoretical analysis. In Witches’ Brew, the attack success decreases from 34–71% to 3–10%; against Sleeper Agent, the attack success decreases from 19–40% to 7–19% [^1] Note that the success rate of typical backdoor attacks is $\sim$90%; thus, the success below 19% means they are ineffective.. Note that to reduce their attack success to $\sim$10%, our defense needs a minimum $\sigma$ of 0.25–0.5. It is also noteworthy that our defense reduces the backdooring success significantly while the certificate does not hold because test-time samples are altered with trigger patterns.
@@ 4.2 Improving Model Utility
A shortcoming of all certified defenses (Levine & Feizi, 2020; Wang et al., 2022; Zhang et al., 2022) is that the certified accuracy is substantially lower than the undefended model’s accuracy. We also observe in the previous subsection that our defense, when we train a model from scratch on the data, denoised with large $\sigma$ values, is not free from the same issue, which hinders their deployment in practice.
In §3, we show that our defense is agnostic to how we train a model, including how we initialize a model’s parameters. We leverage this property and minimize the utility loss by initializing model parameters in specific ways before we train the model on the denoised data. We test two scenarios: a defender can use in-domain data or out-of-domain data to pre-train a model and use its parameters to initialize. We evaluate them against the three most effective attacks: Bullseye Polytope, Witches’ Brew, and Sleeper Agent.
@@
Our defense against $\ell_{\infty}$ attacks at $\sigma$ (%) Att. Initialization ${}^{\dagger}$0.0 0.1 0.25 0.5 1.0 BP N/A (Baseline) ${}^{(93.5)}$100 ${}^{(93.3)}$0.0 ${}^{(92.7)}$0.0 ${}^{(90.8)}$1.0 ${}^{(87.6)}$0.0 WB ${}^{(92.3)}$65.0 ${}^{(86.5)}$9.0 ${}^{(71.9)}$3.0 ${}^{(46.0)}$9.0 ${}^{(41.3)}$7.0 SA ${}^{(92.4)}$34.0 ${}^{(86.2)}$18.5 ${}^{(73.0)}$7.9 ${}^{(47.2)}$8.2 ${}^{(40.1)}$11.5 BP In-domain (CIFAR10) ${}^{(93.5)}$100 ${}^{(93.3)}$0.0 ${}^{(92.6)}$1.0 ${}^{(\mathbf{90.9})}$1.0 ${}^{(\mathbf{87.7})}$2.0 WB ${}^{(92.3)}$65.0 ${}^{(86.5)}$9.0 ${}^{(\mathbf{72.7})}$5.0 ${}^{(\mathbf{47.2})}$7.0 ${}^{(38.6)}$8.0 SA ${}^{(92.4)}$34.0 ${}^{(86.2)}$18.1 ${}^{(\mathbf{73.4})}$8.5 ${}^{(\mathbf{47.8})}$9.9 ${}^{(36.9)}$11.2 BP ${}^{\dagger}$ut-of-domain (ImageNet-21k) ${}^{(85.2)}$4.0 ${}^{(71.1)}$7.0 ${}^{(66.4)}$5.0 ${}^{(58.6)}$5.0 ${}^{(48.7)}$6.0 WB ${}^{(86.6)}$14.0 ${}^{(84.1)}$4.0 ${}^{(\mathbf{79.0})}$0.0 ${}^{(\mathbf{67.7})}$5.0 ${}^{(\mathbf{54.0})}$6.0 SA ${}^{(92.3)}$35.0 ${}^{(86.1)}$18.0 ${}^{(72.9)}$8.5 ${}^{(47.6)}$10.0 ${}^{(37.4)}$10.0 ${}^{\dagger}$esNet18 in Torchvision library; only the latent space dimension differs.

### v2

@@ 3.3 Prediction and Certification
Now we present algorithms for running a prediction with the classifiers trained on $D_{ptr}$ and certifying the robustness of the model’s prediction on a test sample $x$. Our work extends the algorithms presented by Cohen et al. (2019) to clean-label poisoning settings, as shown in Pseudocode 2.
Prediction. To compute the prediction for a test sample $x$, we make an adaptation of the standard randomized smoothing. We train $m$ classifiers on $m$ different noised-then-denoised training datasets, and return the predictions of these classifiers on the target $x$, with the algorithm DTClassify($\cdot$). We run this algorithm for a sufficient number of times (e.g., over 1000 times). The prediction output is then the majority-voted label from these classifiers.
Certification process exactly follows the standard randomized smoothing. We first count the number of occurrences of the most likely label $\hat{y_{A}}$ compared to any other runner-up label, and from this can derive a radius (for the training set perturbations) on which $x$ is guaranteed to be robust.
Results. We train 10k CIFAR10 models using an efficient training pipeline that trains a single model in $15$ seconds on an A100 GPU (for a total of 1.7 GPU days of compute). These models have a clean accuracy of between $92\%$ and $58\%$ depending on the level of noise $\sigma$ introduced, and for the first time can certify non-trivial robustness to clean-label poisoning attacks. A full curve comparing certified accuracy for all given perturbation radii is shown in Figure 1. Instead of applying a diffusion model, we test the randomized smoothing approach proposed by Weber et al. (2023): adding Gaussian random noise directly to the images and train on these (noisy) images. This achieves a similar level of certified accuracy and robustness as our denoising approach, but at a cost of training for 20×20$\times$20 × longer.
@@ 4 Empirical Evaluation
We empirically evaluate the effectiveness of our certified defense against clean-label poisoning. We adapt the poisoning benchmark developed by Schwarzschild et al. (2021).
Poisoning attacks. We evaluate our defense against seven clean-label poisoning attacks: four targeted poisoning (Poison Frogs (Shafahi et al., 2018), Convex Polytope (Zhu et al., 2019), Bullseye Polytope (Aghakhani et al., 2021), Witches’ Brew (Geiping et al., 2021b)) and three backdoor attacks (Label-consistent Backdoor (Turner et al., 2019), Hidden-Trigger Backdoor (Saha et al., 2020), and Sleeper Agent (Souri et al., 2022)). The first three attacks operate in the transfer-learning scenarios (i.e., they assume that the victim fine-tunes a model from some given initialization), whereas the rest assumes the victim trains a model from-scratch.
Because our guarantee is a special case of the one provided by (Weber et al., 2023), the certification holds for both backdooring and clean-label poisoning in $\ell_{2}$.
We evaluate both the transfer-learning and training from-scratch scenarios. We also evaluate against the attackers with two types of knowledge: white-box and black-box. The white-box attackers know the victim model’s internals, while the black-box attackers do not. More details are in Appendix A.
@@
Table 1: Defense effectiveness in transfer-learning scenarios (CIFAR10). We measure the clean accuracy and attack success of the models trained on the denoised training set. Each cell shows the accuracy in the parentheses and the attack success outside. Note that indicates the runs with $\sigma$=0.0, the same as our baseline that trains models without any denoising.
Our defense against $\ell_{2}$ attacks at $\sigma$ (%) Our defense against $\ell_{\infty}$ attacks at $\sigma$ (%) Poisoning attacks Knowledge 0.0 0.1 0.25 0.5 1.0 0.0 0.1 0.25 0.5 1.0 Poison Frog! WB (93.6)99.0 (93.3)0.0 (91.8)1.0 (84.8)0.0 (79.9)1.0 (93.6)68.8 (93.3)0.0 (92.7)0.0 (90.8)0.0 (87.4)0.0 Convex Polytope (93.7)16.2 (93.2)0.0 (91.7)0.0 (86.6)0.0 (77.0)0.0 (93.7)12.2 (93.3)0.0 (92.7)0.0 (90.8)1.0 (87.5)0.0 Bullseye Polytope (93.5)100 (93.3)4.0 (92.6)0.0 (87.5)0.0 (79.2)1.0 (93.5)100 (93.3)0.0 (92.7)0.0 (90.8)1.0 (87.5)0.0 Label-consistent Backdoor - (93.2)1.0 (93.3)0.0 (92.6)0.0 (90.8)1.0 (87.5)0.0 Hidden Trigger Backdoor - (93.4)7.0 (93.3)0.0 (92.6)0.0 (90.8)0.0 (87.5)0.0 Poison Frog! BB (91.6)10.0 (91.2)0.0 (89.6)0.5 (82.9)0.0 (77.8)2.0 (91.7)2.5 (91.3)0.0 (90.3)0.0 (88.8)0.5 (86.2)1.0 Convex Polytope (91.7)3.0 (91.0)0.0 (89.5)0.0 (84.6)0.5 (73.6)1.0 (91.8)2.5 (91.3)0.0 (90.3)0.0 (88.8)0.5 (86.2)1.0 Bullseye Polytope (91.6)9.0 (91.3)0.0 (90.3)0.0 (85.5)0.0 (76.3)1.0 (91.6)8.0 (91.3)0.0 (90.3)0.0 (88.8)0.5 (86.2)0.5 Label-consistent Backdoor - (91.5)1.0 (91.3)0.0 (90.3)0.0 (88.8)0.0 (86.2)1.5 Hidden Trigger Backdoor - (91.6)4.0 (91.2)1.0 (90.3)1.0 (89.3)1.5 (86.3)1.5
@@ 4.1 Effectiveness of Our Denoising Defense
We first show to what extent our defense makes the seven poisoning attacks ineffective. Our defense is attack-agnostic. A defender does not know whether the training data contains poisoning samples or not; instead, the defender always trains a model on the denoised training set. We measure the trained model’s accuracy and attack success and compare them with the baseline, where a model is trained directly without denoising. We evaluate with different defense strengths: $\sigma\in\{0.1,0.25,0.5,1.0\}$. The attacks are with the perturbation bound of 16-pixels in $\ell_{\infty}$-norm and that of 3.5–6.5 in $\ell_{2}$-norm. We extended many poisoning attacks for supporting $\ell_{2}$ bounds. Both $\ell_{\infty}$ and $\ell_{2}$ bounds we use are comparable to each other. We show the conversion between the two bounds in Appendix B. Due to the page limit, the results from Tiny ImageNet are in Appendix D.
@@
Transfer-learning scenarios. Table 1 summarizes our results. Our defense is also effective in mitigating the attacks in transfer-learning. Each cell contains the accuracy and the attack success rate of a model trained on the denoised training set. All the numbers are the averaged values over 100 attacks. Against the white-box attacks, we reduce the success rate of the targeted poisoning attacks to 0–4% at $\sigma$ of 0.1. In the black-box attacks, the targeted poisoning is less effective (with 3–10% success), but still, the defense reduces the success to 0–1%. We observe that the two clean-label backdoor attacks show 1–7% success regardless of the attacker’s knowledge. We thus exclude these attacks from the subsequent evaluation. Using $\sigma$ greater than 0.5 increases the attack success by 0.5–2%. But, this may not because of the successful poisoning attacks but due to the significant decrease in a model’s accuracy (3-9%). Increasing $\sigma$ may defeat a stronger attack but significantly decrease a model’s accuracy. This result is interesting, as in our theoretical analysis shown in § 3, our defense needs $\sigma$ of 0.25, while 0.1 is sufficient to defeat both the attacks.
@@
Training from-scratch scenarios. Table 2 shows our results with the same format as in Table 1. We demonstrate that our defense mitigates both attacks at $\sigma=$ 0.25, aligned to what we’ve seen in our theoretical analysis. In Witches’ Brew, the attack success decreases from 34–71% to 3–10%; against Sleeper Agent, the attack success decreases from 19–40% to 7–19% Note that the success rate of typical backdoor attacks is $\sim$90%; thus, the success below 19% means they are ineffective.Note that to reduce their attack success to $\sim$10%, our defense needs a minimum $\sigma$ of 0.25–0.5. It is also noteworthy that our defense reduces the backdooring success significantly while the certificate does not hold because test-time samples are altered with trigger patterns.
Table 2: Defense effectiveness in training from-scratch scenarios (CIFAR10). We measure the accuracy and attack success of the models trained on the denoised training set. Each cell shows the accuracy in the parentheses and the attack success outside. Note that indicates the runs with $\sigma$=0.0, the same as our baseline that trains models without any denoising. We use an ensemble of four models, and WB and BB stand for the white-box and the black-box attacks, respectively.
Our defense against $\ell_{2}$ attacks at $\sigma$ (%) Our defense against $\ell_{\infty}$ attacks at $\sigma$ (%) Poisoning attacks Knowledge 0.0 0.1 0.25 0.5 1.0 0.0 0.1 0.25 0.5 1.0 Witches’ Brew WB (92.2)71.0 (86.4)54.0 (72.3)10.0 (46.7)11.0 (42.5)10.0 (92.3)65.0 (86.5)9.0 (71.9)3.0 (46.0)9.0 (41.3)7.0 Sleeper Agent Backdoor (92.4)40.5 (84.4)66.5 (71.8)19.5 (46.8)13.0 (39.9)10.5 (92.4)35.0 (86.1)17.0 (73.0)8.0 (47.0)9.5 (39.7)10.0 Witches’ Brew BB (90.1)45.5 (85.9)28.0 (75.5)4.0 (58.8)7.0 (49.0)10.0 (90.0)33.5 (85.8)3.5 (75.5)2.5 (58.8)6.0 (48.7)6.5 Sleeper Agent Backdoor (90.0)39.5 (85.0)44.5 (75.1)14.5 (58.6)9.5 (49.0)8.0 (90.0)18.5 (85.6)11.5 (75.5)7.0 (58.8)8.0 (48.4)8.0
@@ 4.2 Improving Model Utility
A shortcoming of certified defenses (Levine & Feizi, 2020; Wang et al., 2022; Zhang et al., 2022) is that the robust model’s accuracy is substantially lower than that of the undefended model. We also observe in §4.1 that our defense, when we train a model from scratch on the data, denoised with large $\sigma$ values, is not free from the same issue, which hinders their deployment in practice.
In §3, we show that our defense is agnostic to how we train a model, including how we initialize a model’s parameters. We leverage this property and minimize the utility loss by initializing model parameters in specific ways before we train the model. We test two scenarios: a defender can use in-domain data or out-of-domain data to pre-train a model and use its parameters to initialize. We evaluate them against the three most effective attacks shown in Table 1.
@@
Our defense against $\ell_{\infty}$ attacks at $\sigma$ (%) Att. Initialization 0.0 0.1 0.25 0.5 1.0 BP N/A (Baseline) (93.5)100 (93.3)0.0 (92.7)0.0 (90.8)1.0 (87.6)0.0 WB (92.3)65.0 (86.5)9.0 (71.9)3.0 (46.0)9.0 (41.3)7.0 SA (92.4)34.0 (86.2)18.5 (73.0)7.9 (47.2)8.2 (40.1)11.5 BP In-domain (CIFAR10) (93.5)100 (93.3)0.0 (92.6)1.0 (90.9)1.0 (87.7)2.0 WB (92.3)65.0 (86.5)9.0 (72.7)5.0 (47.2)7.0 (38.6)8.0 SA (92.4)34.0 (86.2)18.1 (73.4)8.5 (47.8)9.9 (36.9)11.2 BP Out-of-domain (ImageNet-21k) (85.2)4.0 (71.1)7.0 (66.4)5.0 (58.6)5.0 (48.7)6.0 WB (86.6)14.0 (84.1)4.0 (79.0)0.0 (67.7)5.0 (54.0)6.0 SA (92.3)35.0 (86.1)18.0 (72.9)8.5 (47.6)10.0 (37.4)10.0 ResNet18 in Torchvision library; only the latent space dimension differs.